OpenStudy (anonymous):

help need 5 awnsers fast ill give u medals 2

OpenStudy (anonymous):

Which expression is equivalent to this one? –5(6 + 11) A. (–5) · 6 + 11 B. (–5) + 6 + 11 C. (–5) · 6 · 11 D. (–5) · 6 + (–5) · 11

OpenStudy (sidsiddhartha):

a(b+c)=(a)*(b)+(a)*(c) so which is correct?

OpenStudy (anonymous):

d?????

OpenStudy (sidsiddhartha):

yeah correct

OpenStudy (anonymous):

4 more

OpenStudy (anonymous):

Rewrite the expression without using grouping symbols. 5(x – 19) A. 5 + x – 5 + 19 B. 5x – 519 C. 5x – 5 • 19 D. 5x – 19

OpenStudy (sidsiddhartha):

same property a(b-c)=(a)*(b)-(a)*(c)

OpenStudy (anonymous):

I kinda think its d 2 ???

OpenStudy (sidsiddhartha):

its C 5x – 5 • 19 because u have to multiply 5 with both of them

OpenStudy (anonymous):

ok I have 3 more

OpenStudy (anonymous):

Rewrite the expression by using grouping symbols. y • 4 + y • z A. 4(y + z) B. y (4 + z) C. y • (4 + y) • z D. z(y • 4)

OpenStudy (anonymous):

so is this A I think

OpenStudy (tayy_lovee):

D

OpenStudy (anonymous):

2 more

OpenStudy (anonymous):

What is the solution to the equation? 5(8 + 7) = 5 • x + 5 • 7 A. x = 1 B. x = 5 C. x = 7 D. x = 8

OpenStudy (anonymous):

I think its b

OpenStudy (tayy_lovee):

C